### v1

# Efficient Attention via Pre-Scoring: Prioritizing Informative Keys in Transformers

Abstract
Recent advances in transformer architectures deeply enhance long-context language modeling. Among them, HyperAttention achieves competitive efficiency by combining a single-level LSH-based clustering with uniform residual sampling. However, HyperAttention fails to find all significant keys, which in turn raises the overall perplexity. We propose a pre-scoring mechanism to assist HyperAttention to prioritize significant keys. We introduce three scoring methods: $k$-means and kernel $k$-means clustering, $k$-median clustering, and leverage score-based ranking (inspired by LevAttention) to filter keys effectively. We further replace HyperAttention’s original uniform residual sampling entirely, relying exclusively on our pre-scoring mechanism. Experiments on ChatGLM2 (131k token context) reduce perplexity from 12 to 8.3, which outperforms standard HyperAttention. Moreover, when running on the Vision-Transformer (ViT), our method shows that it can guarantee similar accuracy compared with LevAttention, and will surpass LevAttention given specific parameters. Although this method introduces some computational overhead, its combination with HyperAttention remains 20 times faster than FlashAttention, providing a balanced trade-off between speed and modeling accuracy. Our results highlight the effectiveness of integrating pre-scoring into hierarchical attention mechanisms, significantly improving transformer efficiency.
1 Introduction
Transformer-based large language models (LLMs) now underpin cutting-edge performance across diverse applications, including computer vision [4] and text classification [13], yet their quadratic self-attention cost remains a persistent barrier to efficient long-context processing. Especially for Transformer-based attention mechanisms, the time complexity rises quadratically with respect to the sequence length. Without effective strategies to alleviate this cost, deploying transformer models on tasks requiring truly long contexts remains impractical, limiting both training scalability and inference efficiency.
Consider the input of each attention layer, which is typically represented as an $n\times d$ matrix $X$, where $n$ denotes the context length and $d$ is the embedding dimension of the tokens. From this input matrix, we generate three distinct matrices by applying learned linear transformations. Specifically, we compute the query matrix $Q=X\cdot W_{Q}$, the key matrix $K=X\cdot W_{K}$, and the value matrix $V=X\cdot W_{V}$, where $W_{Q}$, $W_{K}$, and $W_{V}$ are learned parameter matrices of dimensions $d\times d$. Various methods have been developed to improve efficiency when processing $Q$, $K$, and $V$. For example, Performer [5] replaces the standard softmax attention with kernel methods. This remarkably reduces computational complexity, yet it involves potential kernel approximation-induced errors. FlashAttention [6] facilitates computation by optimizing memory access patterns during the matrix computations. Despite its advantages, specialized hardware support is required for optimal performance.
In addition, HyperAttention [9] reduces the computational burden using approximate matrix product and locality sensitive hashing. These methods improve efficiency at the cost of some degradation in perplexity. LevAttention [10] selects a fixed set of keys independent of the queries to try to capture all heavy attention scores. However, it is less effective in capturing query-specific attention patterns.
Inspired by HyperAttention and LevAttention, we propose a pre-scoring mechanism. Namely, we use clustering-based filtering to prioritize informative keys before applying HyperAttention.
Our Results We first show that clustering-based pre-scoring consistently outperforms leverage-score selection: for example, on a Vision Transformer (ViT) Large model [8] (85.85% baseline accuracy on ImageNet-1k), sampling 128 keys via K-means retains 84.46% accuracy, compared to only 77.17% when using leverage scores (top-128). Furthermore, by integrating our clustering-based pre-scoring into HyperAttention, we reduce ChatGLM2-6b-32k’s perplexity on the Longbench data set [3] from 12 to 8.3, outperforming both the original HyperAttention and the variant augmented with leverage-score selection (LevAttention+HyperAttention) [14, 10]. Moreover, our method preserves the computational efficiency advantages of HyperAttention over standard self-attention and FlashAttention [6].
In addition to our empirical results, we revisit the planted-subspace model introduced for LevAttention (Section 4.4 of [10]), and we show that our pre-scoring clustering-based methods recover the same planted model guarantees of LevAttention, in terms of recovering all heavy keys for a query, with high probability. Thus, theoretically our methods are no worse than those of LevAttention in this natural planted model. We further introduce an alternative planted-subspace model in which K-means clustering provably recovers keys with large leverage score, showing again that K-means pre-scoring is as powerful as leverage score pre-scoring for natural models. This may help explain why our empirical results consistently outperform those of LevAttention [10].
2 Preliminaries
We define the attention mechanism $Att=D^{-1}AV$. Here, $A$ is defined as $A:=\exp(QK^{\top})$, while $D$ is a diagonal matrix with $D_{i,i}=\|A_{i,:}\|_{1}$ for each $i\in[n]=\{1,2,\ldots,n\}$. We refer to the matrix $A$ as the attention matrix. Explicitly calculating the attention matrix requires $\Theta(n^{2}d)$ time and $\Theta(n^{2})$ memory, which can be prohibitive.
Han et al. introduce HyperAttention [9], which addresses the quadratic bottleneck of vanilla self-attention by hashing queries and keys with an angular locality-sensitive hash (LSH) function and then ordering buckets so that adjacent buckets have small Hamming distance. Scores are computed only for pairs that fall into the same hash bucket, significantly reducing the time and memory cost on typical data distributions. Their method performs an additional randomized low-rank compression inside each bucket, further reducing constant factors. Its main limitation is data independence: because bucket membership is determined solely by the hash function, weak but semantically crucial long-range links may never collide and can therefore be missed.
LevAttention [10] takes a complementary view. It first sketches the key matrix in $O(nd)$ time to approximate statistical leverage scores for each row of the key matrix, then forms a Universal set $U=\{i:\operatorname{LS}(K_{i})\geq\epsilon\}$ that, for polynomial-based attention (rather than the standard softmax attention), is guaranteed to contain all attention scores whose weight exceeds a user-chosen threshold $\epsilon>0$. Looking only at $U$ therefore achieves perfect recall of heavy attention scores, independent of positional locality. However, when $\epsilon$ is very small, $|U|$ can be large as $n$, eliminating any savings.
Moreover, a uniform evaluation within the set $U$ wastes time in many pairs with low impact. By first constructing $U$ to ensure theoretical coverage and then applying HyperAttention’s locality-sensitive hashing inside this recall-guaranteed set, our hybrid approach combines the strengths of both methods: it retains subquadratic complexity and provably captures every $\epsilon$-heavy attention.
Our goal is to accelerate approximation methods for transformers while maintaining high accuracy. Rather than using the universal set $U$ of Kannan et al. [10], we suggest other sets based on clustering methods to guide approximation algorithms such as HyperAttention [9] to find large attention scores.
3 Algorithm
A central motivation for our pre-scoring approach is the hypothesis that computationally efficient methods, such as clustering, can effectively identify and prioritize the most salient keys within the key matrix $K\in\mathbb{R}^{n\times d_{k}}$ (where $n$ is sequence length and $d_{k}$ is the key dimension). This idea is supported by Axiotis et al. [2]. They developed a cluster-based sensitivity sampling method to enhance data selection efficiency for large-scale model training. This principle resonates with techniques such as LevAttention [10], which use statistical leverage scores to sample influential data points. To provide rigorous grounding for clustering, and specifically K-means or K-median) as a pre-scoring mechanism, we analyze its performance and relationship to leverage scores under a structured data model.
3.1 Pre-Scored HyperAttention
Algorithm 1 ranks the $n$ keys in a single pass. Given a stochastically perturbed matrix $K^{\prime}$ of keys, it offers two routes: (i) a one-shot $k$-means/$k$-median call that returns the $s$ closest keys to $k=d{+}1$ centroids, or (ii) a fast leverage-score sketch. Theorems 1 and 2 in Section 3 guarantee that both routes isolate all $\Theta(\epsilon)$-heavy keys with exponentially small failure probability. The full model and assumptions are given in Section § 3).
Algorithm 1 PreScore: Rank Keys via Clustering or Leverage
[…]
2:$K^{\prime}\leftarrow K+\mathcal{N}(0,\sigma^{2}I_{d_{k}})$triangleright optional noise
3:if method == KMEANS or KMEDIAN then
4:     $\{C_{j},\mu_{j}\}_{j=1}^{k}\leftarrow\textsc{KMeans}(K^{\prime},k)$
5:     $S\leftarrow$ indices of the $s$ keys nearest to their centroids
6$\triangleright$ LEVERAGE branch
7:     $h\leftarrow\textsc{ApproxLeverage}(K^{\prime})$triangleright $O(nd_{k}\!\log d_{k})$ time
8:     $S\leftarrow$ top-$s$ indices by $h$
9:end if
10:return index set $S$
Lines 3–7 execute in $O(nd_{k})$ time for clustering, and $O(nd_{k}\!\log d_{k})$ for leverage scores.
Algorithm 2 wraps the pre-scoring routine around HyperAttention. It first calls PreScore to obtain the retained key index set $S$; if fewer than $\delta n$ keys survive, it falls back to vanilla HyperAttention to match the baseline’s worst-case runtime. Otherwise, it applies HyperAttention only to the $|S|=s$ scored keys, yielding a near-linear $O(nd_{k}+sd_{k})$ layer cost.
Algorithm 2 Pre-Scored HyperAttention
[…]
2:$S\leftarrow\textsc{PreScore}(K,k,s,\sigma,\textit{method})$triangleright Algorithm 1
3:if $|S|<\delta n$ then $\triangleright$ robust fallback
4:     return $\textsc{HyperAttention}(Q,K,V)$
5:end if
6:$\textit{Att}_{\text{out}}\leftarrow\textsc{HyperAttention}(Q,K[S],V[S])$
7:return $\textit{Att}_{\text{out}}$
3.2 Structural Guarantees
Matrix Structure and Assumptions:
We give a new planted model for which cluster-based prescoring recovers all large leverage scores.
Let $A\in\mathbb{R}^{n\times d}$ be a matrix generated as follows:
  1. 1.
    There are $d$ disjoint sets of row indices, $S_{1},\dots,S_{d}$, each of size $m=\lceil 1/\epsilon\rceil$ for some small $\epsilon\in(0,1)$.
  2. 2.
    Let $S_{0}=\{1,\dots,n\}\setminus\bigcup_{j=1}^{d}S_{j}$ be the set of remaining row indices. We assume $n\gg dm$, so $|S_{0}|=n(1-o(1))$.
  3. 3.
    Let $v_{1},\dots,v_{d}\in\mathbb{R}^{d}$ be an orthonormal basis for $\mathbb{R}^{d}$.
  4. 4.
    For each $j\in\{1,\dots,d\}$, and for each row index $i\in S_{j}$, the $i$-th row of $A$ is given by $A_{i}=v_{j}+\delta_{i,j}$, where $\delta_{i,j}\sim\mathcal{N}(0,\sigma_{S}^{2}I_{d})$ are independent and identically distributed (i.i.d.) Gaussian noise vectors.
  5. 5.
    For each row index $i\in S_{0}$, the $i$-th row of $A$ is given by $A_{i}=\eta_{i}$, where $\eta_{i}\sim\mathcal{N}(0,\sigma_{N}^{2}I_{d})$ are i.i.d. Gaussian noise vectors.
  6. 6.
    The noise scales satisfy $\sigma_{S}^{2}=c_{S}/d$ and $\sigma_{N}^{2}=c_{N}/(n\epsilon)$ for sufficiently small positive constants $c_{S},c_{N}$. This implies that the noise variance within groups $d\sigma_{S}^{2}=c_{S}$ and within the noise group $d\sigma_{N}^{2}=dc_{N}/(n\epsilon)$ are small.
  7. 7.
    Row-norm regularity: $\|A_{i}\|_{2}=1$ for all $i$.
Remark
The correlation bounds ($P_{1}$)–($P_{2}$) do not prevent a few rows from having arbitrarily large $\ell_{2}$ norm. Appendix 7 presents a simple construction where $\delta_{1}=\delta_{2}=0$ yet the optimal $k$-means mis-clusters because some high-norm noise rows dominate the objective. In the large-norm construction, despite perfect orthogonality ($\delta_{1}=\delta_{2}=0$), the $k$-means objective is overwhelmed by the $M^{2}$-scaled distances of the noise rows. As a result, clusters are “stolen” from the true signal set $S$ in order to mitigate the huge distortions induced by those high-norm points, and $S$ is not recovered. This shows unambiguously that without enforcing $\|A_{i}\|_{2}=1$, clustering can fail completely. To rule this out, we henceforth assume all rows are $\ell_{2}$-normalized (equivalently, have comparable variance), i.e. $\|A_{i}\|_{2}=1$ for every $i$. This “norm-regularized planted-subspace model” is what the following analysis relies on.
Recall that for any row $i$, its leverage score is $h_{i}=A_{i}\,(A^{\top}A)^{-1}A_{i}^{\top}=\sup_{\|x\|=1}\frac{(A_{i}^{\top}x)^{2}}{\|Ax\|^{2}}.$ Based on our assumption, we have the following:
Lemma 1 (Upper Bound on Noise Leverage).
In this model, for each row $i\in S_{0}$ we have $h_{i}\leq\frac{\|A_{i}\|^{2}}{\sigma_{\min}^{2}}=O\bigl{(}\tfrac{d}{n}\bigr{)},$ where $\sigma_{\min}^{2}=\lambda_{\min}(A^{\top}A)=\Theta(1)$.
Lemma 2 (Lower Bound on Signal Leverage).
For each row $i\in S_{j}$, choose unit $x=v_{j}$. Then $h_{i}\geq\frac{(A_{i}^{\top}v_{j})^{2}}{\|Av_{j}\|^{2}}=\Theta(\epsilon).$
The above lemmas are standard; we refer to the supplementary for their proofs. Letting $A$ be the key matrix $K$, we have: $h_{j}=\sup_{\|x\|=1}\frac{(k_{j}^{\top}x)^{2}}{\|Kx\|^{2}},$ and $\|Kx\|^{2}\geq\sigma_{\min}^{2}=\Theta(1)$.
Theorem 1 (Leverage-Score Separation).
Under the same model, with high probability
$\forall\,i\in S_{j}:\;h_{i}=\Theta(\epsilon),\quad\forall\,i\in S_{0}:\;h_{i}=O\bigl{(}\tfrac{d}{n}\bigr{)}.$
Given Lemma 1 and Lemma 2, we immediately obtain Theorem 1 as follows. Lemma 1 gives, for any noise row $i\in S_{0}$, $h_{i}\leq O\bigl{(}\tfrac{d}{n}\bigr{)}$ and from Lemma 2, $h_{i}\geq\Theta(\epsilon).$ Since by assumption $n\gg d/\epsilon$ we have $O(d/n)=o(\epsilon)$. Hence with high probability
$\forall\,i\in S_{j}:\;h_{i}=\Theta(\epsilon),\quad\forall\,i\in S_{0}:\;h_{i}=O\bigl{(}\tfrac{d}{n}\bigr{)}.$
A formal proof of Theorem 1 is in the supplementary.
Theorem 2 (K-means Clustering).
Under the same assumptions as Theorem 1, and assuming $c_{S}$ and $c_{N}$ are sufficiently small (e.g., $c_{S}<1/2$ and $dc_{N}/(n\epsilon)<1/2$), with probability at least $1-\exp(-\Omega(\min(m,n-dm,d)))$), the $k$-means algorithm with $k=d+1$ applied to the rows of $A$ converges to a clustering where:
  1. 1.
    There are $d$ clusters, $C_{1},\dots,C_{d}$, such that for each $j\in\{1,\dots,d\}$, all rows in $S_{j}$ are assigned to cluster $C_{j}$. The centroid $\mu_{j}$ of $C_{j}$ satisfies $\|\mu_{j}-v_{j}\|=O(\sigma_{S}/\sqrt{m})$.
  2. 2.
    There is one cluster $C_{0}$ containing all rows from $S_{0}$. The centroid $\mu_{0}$ of $C_{0}$ satisfies $\|\mu_{0}-\mathbf{0}\|=O(\sigma_{N}/\sqrt{n-dm})$.
Running K-means with $k=d+1$ clusters on $\{k_{j}\}$ yields centroids $\mu_{1},\dots,\mu_{d},\mu_{d+1}$. Under the above separations, an optimal solution aligns with $C_{i}=G_{i}$ for $i=1,\dots,d$ and $C_{d+1}=S$, since $\sum_{k_{j}\in G_{i}}\|k_{j}-u_{i}\|^{2}=0,\quad\sum_{s\in S}\|s-\mu_{d+1}\|^{2}\leq\sum_{s\in S}\|s-\mu\|^{2}\quad\forall\mu$, and any deviation incurs a large additional within-cluster sum of squares (WSS). In fact, by Theorem 1 the true partition $C_{j}=S_{j}$ ($j=1,\dots,d$) and $C_{d+1}=S_{0}$ has total within-cluster cost
$\sum_{j=1}^{d}\sum_{i\in S_{j}}\|A_{i}-v_{j}\|^{2}+\sum_{i\in S_{0}}\|A_{i}\|^{2}=O(m\,\sigma_{S}^{2})+O((n-dm)\,\sigma_{N}^{2})=o(1)\,,$
whereas moving any single point to the wrong cluster incurs an additional penalty of at least $\min\{\|v_{j}-v_{k}\|^{2},\|v_{j}\|^{2}\}-o(1)=1-o(1)>0.$ Hence any mis-assignment increases the total sum-of-squares, making the true grouping the unique global minimizer of the $k$-means objective. Thus, clustering identifies those rows with $h_{j}\geq\epsilon$, matching LevAttention’s heavy-key selection.
Corollary 1 (Singleton case of Theorem 2).
Setting $m=1$ (so $\epsilon=1$ in this special case) in Theorem 2 shows that, with the same probability, the optimal $k$-means clustering with $k=d+1$ places every signal row in its own cluster and gathers all noise rows in $C_{0}$.
Proof Sketch.
With $m=1$, each heavy row $K_{j}\in S$ contributes zero within-cluster distortion when isolated. If instead $K_{j}$ is merged with any other row $K_{x}$, the centroid error lower bound
$\|K_{j}-\mu\|_{2}^{2}\;\geq\;\tfrac{1}{2}\|K_{j}-K_{x}\|_{2}^{2}\;\geq\;\tfrac{1}{2}D_{\min}$
applies, where $D_{\min}>0$ is the minimum inter-point separation guaranteed by Theorem 1. Since $D_{\min}/2$ remains a positive constant, any such mis-assignment strictly increases the total $k$-Means cost. Finally, choosing $k=d+1$ reserves one centroid per heavy key and one for all noise rows, matching the leverage-score separation. For a detailed proof, see Appendix 9
Connection to the planted model and the constant gap $D_{\min}$.
The Gaussian planted model studied in Section 4.4 of [10] draws signal and noise rows from two covariance profiles whose variances differ by a fixed ratio. Standard concentration shows that this forces a constant lower bound
$D_{\min}=2(1-\vartheta_{1})\!\geq\!\tfrac{3}{2}$
on the squared Euclidean distance between any signal and any noise row whenever the variance-ratio parameter satisfies $\vartheta_{1}<\tfrac{1}{4}$. That same constant gap is exactly what drives Corollary 1: putting a signal row into a mixed cluster would increase the $k$-means objective by at least $D_{\min}/2$, so the optimal solution (with $k=d+1$) must isolate every signal row and pool all noise rows in $C_{0}$. Thereby, the planted model supplies a probabilistic guarantee for the deterministic separation our corollary needs, confirming that the singleton clustering phenomenon emerges naturally whenever the signal-to-noise variance ratio is sufficiently small.
In addition to our theorems above, this framework extends naturally to any $\ell_{p}$ norm where $p>0$, allowing us to generalize our pre-scoring mechanism. Consider a key matrix $K=\{k_{j}\}_{j=1}^{n}\subset\mathbb{R}^{d_{k}}$ drawn from a mixture of $d$ well-separated “heavy” centers and a bulk of “light” points, as in our structured data model. We define the $\ell_{p}$-sensitivity of each key $k_{j}$ following [12], which quantifies the importance of keys under the $\ell_{p}$ norm.
To adapt our clustering approach, we employ Minkowski-$k$-means, which minimizes the following: $\sum_{j=1}^{n}\min_{i\in[d]}\|k_{j}-\mu_{i}\|_{p}^{p}.$ This method clusters keys by minimizing the $p$-th power of their $\ell_{p}$ distances to the nearest centroid. Under the same separation conditions as in Lemma 2, Minkowski-$k$-means[11] accurately recovers the top-heavy keys according to their $\ell_{p}$-sensitivity. Specifically, we show:
Claim 1 ($\ell_{p}$-Generalization).
Under the structured model of §3.3 (with disjoint signal sets $S_{1},\dots,S_{d}$, noise set $S_{0}$, and $A_{i}=v_{j}+\delta_{i,j}$ for $i\in S_{j}$, $A_{i}=\eta_{i}$ for $i\in S_{0}$, and noise scales $\sigma_{S}^{2}=c_{S}/d$, $\sigma_{N}^{2}=c_{N}/(n\epsilon)$), running $k$-means with the $\ell_{p}$ metric (i.e., using distances $\|x-y\|_{p}^{p}$) and $k=d+1$ recovers exactly the true clusters $S_{1},\dots,S_{d},S_{0}$, provided $c_{S}$ and $c_{N}$ are sufficiently small.
For a brief proof of this claim, we replace every squared-norm ($p=2$) in the analysis of lemma 1 with the $p$-th power norm, and the separation conditions ensure that the heavy centers remain distinguishable. Based on our proof in 8.5, we conclude: In the $\ell_{p}^{p}$ metric one checks that $\|v_{j}-0\|_{p}^{p}=1$ and $\|v_{j}-v_{k}\|^{2}=2$ for $j\neq k$ so the minimum inter-centroid separation is $\Delta_{\min}=1$, while standard bounds show that each point’s $p$-th power deviation $\delta_{\max}=O(c_{S}^{p/2}d^{1-p/2})=o(1)$. Hence any mis-assignment raises the $k$-means cost by at least $\Delta_{\min}-2\delta_{\max}>0$, making the true partition the unique global minimizer. Since $\delta_{\max}=o(1)$ holds with probability at least $1-\exp(-\Omega(\min(m,n-dm)))$, $\ell_{p}$-$k$-means recovers the true clusters with high probability.
This generalization enables our method to prioritize informative keys under various $\ell_{p}$ metrics, which may be advantageous for different data distributions or model architectures. By restricting the queries to attend only to the sampled keys selected by the method above, we ensure focused attention mechanisms while preserving the framework’s adaptability. We provide both deterministic and probabilistic analyses of this connection in Appendix 8.
4 Experiments and Results
To evaluate our proposed pre-scoring attention algorithm, we conducted comprehensive experiments assessing runtime efficiency, perplexity performance, and applicability to Vision Transformers (ViTs) [8]. We compared three variants—K-means+Hyper, K-median+Hyper, and Lev+Hyper—against baseline HyperAttention [9] and FlashAttention [6]. We also analyzed the key clustering performance of K-means relative to ViT’s standard self-attention. All the experiments were done on a single NVIDIA A100 GPU with 40 GB memory or a single NVIDIA L4 GPU with 24 GB memory.
4.1 Speed Comparison on Flash Attention
Figure 1: Single layer speed test by only forward pass
Figure 2: Single layer speed test by forward and backward pass
We test the speed-up factor for each layer compared to that of Flash Attention. This speedup factor is the ratio of Flash Attention’s runtime to that of each tested method on a per-layer basis. From the results, all combinations, Lev + Hyper, K-means + Hyper and K-median + Hyper, outperform Flash Attention for sufficiently long sequences, demonstrating the advantage of HyperAttention-based methods. Compared to the original HyperAttention, these methods can generate a mild acceleration, with performance becoming more remarkable starting at $2^{13}$ with a speedup factor of around $3$ to $4$ in Figure 2. Such tradeoffs are from the time complexity of the pre-scoring algorithm in forward selection parts, which is $O(N\cdot d^{2})$ for Lev+Hyper and $O(N\cdot d\cdot k)$ for K-means/median+Hyper, where $k$ is the number of clusters in the K-means/median part. Because of the huge size of the key matrix, having $d\gg k$ is a common situation. In this case, it is also reasonable to conclude that the additional complexity is roughly $O(N\cdot d)$, a near-linear complexity dependent solely on the dimensions of $K$. Comparing these two options, Lev+Hyper exhibits the best scalability, closely matching HyperAttention on longer sequences, while K-median+Hyper requires slightly higher computational cost due to its clustering complexity. Our pre-scoring overhead is most pronounced in the forward pass, as the backward pass adheres to HyperAttention’s standard pipeline. As a result, extending pre-scoring to both passes could potentially narrow the speedup factor.
4.2 Accuracy Comparison
While their speedup factors grow more gradually, our HyperAttention-based methods offer a balanced trade-off between runtime and perplexity, as observed in the speed comparison. Similarly to HyperAttention’s test [9], we use the Longbench dataset [3] on ChatGLM2-6b-32k model [14]. To compare the efficacy of different methods under full-layer replacement, we incorporate K-means/K-median and LevAttention with our scoring mechanism before sending the scored data to the HyperAttention algorithm.
Figure 3: The perplexity performance of various key selection strategies—K-means, K-median, and Leverage Score (Lev)—under different values of $k$ (the number of selected keys), where each $k$ is sampled as a power of $2$ (e.g., $128$, $2048$, $8192$)
Figure 3 shows the methods that incorporate sampling: K-means, K-median, and Lev, consistently outperform standard HyperAttention across various top-$k$ values. These methods exhibit a U-shaped trend: The perplexity initially decreases as $k$ increases, reaching an optimal range around $2048$ to $8192$. Note that the original HyperAttention has perplexity=12, our pre-scoring algorithm will lower the perplexity for some top-k greater than threshold values between $128$ and $2048$. With this information and fix a speedup module, if we can calculate this value $\tau$ as a threshold to activate the residual sampling module, our algorithm will always guarantee lower perplexities than HyperAttention for $top_k|>\tau$. At top_k=2048, we obtained the highest accuracy that perplexity=10.38, which is a 12.5% improvement. In addition, setting min_seq_len >= n_query ensures that the model bypasses fallback mechanisms in the causal attention branch, enabling full use of blockwise optimization even at shorter sequence lengths. The perplexity reaches a minimum of $8.31$ under the condition, representing a substantial improvement over the baseline perplexity of $12$ with a percentage improvement of 30.8% . Moreover, an excessive increase in top-k, such as k=16384, will result in a rise in perplexity, and degenerate back to the original HyperAttention results[9].
We further tested Gaussian kernel k-means, which computes distance to centroids via the kernel method. The test results are shown in the table below. This method has its best representation with perplexity=10.06 at top-k=8192 when Sample Size=256. Similar to previous methods, the perplexity shown by this reaches the lowest value at top-k=0 when Sample Size=0.
Table 1: PPL comparison for Gaussian Kernel K-means across configurations.
Top K Sample Size PPL
0 256 17.5410
32 256 18.5715
128 256 19.6897
512 256 15.6162
1024 256 14.3289
2048 256 11.8682
4096 256 10.5833
8192 256 10.0670
16384 256 11.8495
0 0 10.4122
32 0 10.7094
128 0 11.6935
512 0 11.2801
1024 0 11.8827
2048 0 13.6853
4096 0 17.6743
8192 0 22.8829
16384 0 34.6969
4.3 Monkey Patching Vision-Transformer
We replaced the standard softmax-based self-attention layers in the Vision Transformer (ViT) [8] with our proposed K-means sampling attention mechanism by letting queries $Q$ only attend to a subset $S$ of keys $K$ chosen by our algorithm 1. We also mask the value matrix $V$ with our subset $S$ to align with the original output shape. For the baseline, we used the pretrained vit_small_patch16_224 and vit_large_patch16_224 models, which achieved top-1 accuracies of 85.11% and 85.85%, respectively, on the ImageNet-1k validation set [7]. Our custom attention mechanism replaces the full attention computation with a K-means clustering approach that samples a subset of key vectors per head. We varied the number of clusters and sampled keys to evaluate the performance. In the ViT-Small variant, when the number of clusters was fixed at 4, reducing the number of sampled keys to 32 resulted in a drastic accuracy drop (31.34%), while increasing the sample count to 64, 96, and 128 gradually improved the accuracy to 61.31%, 74.21%, and 80.05%, respectively. A further increase to 6 clusters with 128 sampled keys yielded a marginal improvement (80.49%). In contrast, the ViT-Large variant exhibited higher robustness: with 4 clusters, the accuracy improved from 53.05% with 32 samples to 78.10%, 82.89%, and 84.46% for 64, 96, and 128 samples, respectively, while increasing the cluster count to 6 with 128 sampled keys maintained an accuracy of 84.46%. These results, summarized in Table 2., indicate that our K-means sampling attention can closely approximate the performance of the full self-attention mechanism if a sufficient number of key vectors is sampled. Our method achieve better performance than LevAttention [10] with pretrained ViT models and share similar accuracy trade-offs to LevAttention with models trained from scratch using their updated leverage-score based attention mechanism. More detailed results of LevAttention are included in Appendix 10.
Table 2: Accuracy Results for Monkey Patched ViT Models using K-means Prescoring
Model Configuration Accuracy
S/16 Base model 85.11%
num_cluster: 4, num_sample: 32 31.34%
num_cluster: 4, num_sample: 64 61.31%
num_cluster: 4, num_sample: 96 74.21%
num_cluster: 4, num_sample: 128 80.05%
num_cluster: 6, num_sample: 128 80.49%
L/16 Base model 85.85%
num_cluster: 4, num_sample: 32 53.05%
num_cluster: 4, num_sample: 64 78.10%
num_cluster: 4, num_sample: 96 82.89%
num_cluster: 4, num_sample: 128 84.46%
num_cluster: 6, num_sample: 128 84.46%
4.4 Heavy Attention Coverage
We evaluated our custom attention mechanism that employs K-means and K-Median sampling to approximate the full attention distribution. This experiment was also conducted using ViT base model with patch 16 [8] on Imagenet-1k validation set [7].We compared the original attention output matrix with the attention output matrix after applying our clustering-based prescoring mechanism for analysis. We vary the number of sampled keys (32, 64, and 128) and adjust the threshold parameter $\epsilon$ (with values 0.01, 0.1, and 0.3) to measure the median percentage of heavy attention entries captured. An entry of attention matrix $A$ is considered heavy if $A_{ij}>\epsilon$ for $0\leq i,j\leq N$. From Figure 5 and Figure 5, the capture percentage increases as $\epsilon$ or the number of keys sampled increases and K-Means has some marginal performance increase compared to K-Median. Additionally, we tested the number of top columns that contain the most heavy attention entries, and how well these can be captured by both sampling approaches. The result is shown in Appendix 10.1. While the overall heavy attention entries coverage rate shows a linear relationship, the top-$k$ coverage remains the same as the configurations change (where $k$ aligns with our number of keys sampled).
[…]
Figure 4: K-Means: Median percentage vs. sampled keys.
[…]
Figure 5: K-Median: Median percentage vs. sampled keys.
5 Conclusion
In this work, we proposed a novel pre-scoring mechanism that integrates clustering-based key selection methods to improve hierarchical attention mechanisms. By selectively prioritizing informative keys, we overcome limitations of HyperAttention’s uniform residual sampling, achieving significant perplexity improvements (from 12 down to 8.3) and computational efficiency advantages over FlashAttention. Empirical results across language models and vision transformers validate our method’s effectiveness and generality. Mathematically, we also used planted-subspace model to prove that our clustering-based scoring is as powerful as leverage score pre-scoring for natural models. Our work motivates the study of a wide variety of scoring mechanisms that go far beyond the initial use of leverage scores in [10]. Future research directions may include:
(1) Investigating adaptive pre-scoring strategies dynamically tailored during training, (2) Extending the pre-scoring paradigm to other Transformer variants beyond language modeling, such as multimodal and cross-modal architectures, and (3) Further optimization of clustering or leverage-scoring algorithms to minimize computational overhead, enhancing applicability in practical scenarios.